A voltaic cell consists of a standard hydrogen electrode in one half-cell and a Cu/Cu2+ half-cell. Calculate [Cu2+] when E cell is 0.160 V.
Answer #1

         H2(g) -------------> 2H^+ (aq) + 2e^-          E0 = 0v

        Cu^2+ (aq) + 2e^- -----------> Cu(s)             E0 = 0.34v

-----------------------------------------------------------------------------------

Cu^2+ (aq) + H2(g) -----------> 2H^+ (aq) + Cu(s)     E0cell = 0.34v

n = 2

Ecell   = 0.16v

Ecell   = E0cell - 0.0592/n log[H^+]^2/[Cu^2+]PH2

0.16 = 0.34 - 0.0592/2 log1/[Cu^2+]*1

0.16    = 0.34 - 0.0296log1/[Cu^2+]

0.16 -0.34   = -0.0296log1/[Cu^2+]

-0.18            = -0.0296log1/[Cu^2+]

log1/[Cu^2+]    = 0.18/0.0296

log1/[Cu^2+]       = 6.08

1/[Cu^2+]            = 10^6.08

1/[Cu^2+]             = 1.2*10^6

[Cu^2+]           =   8.34*10^-7 M   >>>>answer